ROVIBRATIONAL LEVEL STRUCTURE OF ACETYLENE BY STIMULATED EMISSION PUMPING: CORIOLIS COUPLING BETWEEN ν_{2} + 4ν_{4} AND 7ν_{4}

\underline{DAVID B. MOSS}, Division of Science and Mathematics, College of General Studies, Boston University, Boston, MA 02215; RICHARD DUAN, MATTHEW P. JACOBSON, JONATHAN P. O’BRIEN, and ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

Stimulated emission pumping (SEP) spectroscopy has been used to investigate the rovibrational level structure of S_{0} acetylene with 0.1 cm^{−1} resolution at a vibrational energy of approximately 4400 cm^{−1}. Both the e and f parity levels of the expected zero-order bright state ν_{2} + 4ν_{4} ( ℓ = 0, 2 ) are observed to be doubled by a rotationally inhomogeneous perturbation. SEP spectra obtained using S_{1} intermediate levels with K = 1 vs. K = 0 reveal complementary intensity patterns, indicating that the perturbation follows a selection rule for vibrational angular momentum of Δℓ = odd. The perturbation is assigned to a Coriolis interaction between ν_{2} + 4ν_{4} and the pure bending dark state 7ν_{4}. Zero-order vibrational energies and B values, as well as the Coriolis coupling coefficient, are derived from a fit of greater than 60 rovibrational levels of both parities with J = 1 to 13. Parameters derived from the fit are compared with predictions based on previous infrared and dispersed fluorescence studies. Axis-switching effects are also found to contribute significantly to the observed intensities.